Two students, Jim and Joan are registered for the same class and attend independently to each other, Jim 91,9% of the time and Joan 84.3% of the time. What is the probability that on any given day
a) at least on of them will be in class?
b) neither will be in class?

Answer by stanbon(75887)

Two students, Jim and Joan are registered for the same class and attend independently to each other, Jim 91,9% of the time and Joan 84.3% of the time. What is the probability that on any given day
a) at least on of them will be in class?
P(at least one) = 1 - P(neither in class)
= 1 - 0.081*0.157 = 0.9873
--------------------------------------
b) neither will be in class?
P(neither) = 0.081*0.157 = 0.0127
=======================================
cheers,
Stan H.
==============